need answer for the below question
he sun of the ages of a mother and a daughter is 50 years. Also 5 years ago the mother’s age was 7 times the age of the daughter. The present ages of the mother and the daughter respectively are:

Found 2 solutions by Cromlix, rothauserc:
Answer by Cromlix(4381)

The sum of the ages of a mother and a daughter is 50 years.
M + D = 50
M = 50 - D
Also 5 years ago the mother’s age was 7 times the age of the daughter.
M - 5 = 7(D - 5)
Replace M with 50 - D
50 - D - 5 = 7D - 35
8D = 80
D = 10
Currently Daughter is 10 years old
Mother is 40 years old.
Hope this helps.
:-)

Answer by rothauserc(4718)

let x be the age of the mother and y be the age of the daughter
x +y = 50
x-5 = 7(y-5)
x = 7(y-5)+5
x = 7y -35 +5
x = 7y -30
now substitute for x in the first equation
7y -30 +y =50
8y = 80
y = 10
x = 40
The mother is 40 and the daughter is 10